OpenStudy (anonymous):

Given sinu=4/5, with u in quadrant 1, and cosv=7/25, with v in quadrant 1, Find cos(u+v) Here are the choices A) (-3/5) B) (-44/125) C) (117/125) D) (4/5)

jimthompson5910 (jim_thompson5910):

hint: cos(u+v) = cos(u)cos(v) - sin(u)sin(v)

jimthompson5910 (jim_thompson5910):

so you'll need to know the values of cos(u) and sin(v)

OpenStudy (anonymous):

still confused....

jimthompson5910 (jim_thompson5910):

if sin(u) = 4/5, then what is cos(u) ?

OpenStudy (anonymous):

well wouldn't they just be the same?

jimthompson5910 (jim_thompson5910):

no they aren't

OpenStudy (anonymous):

they how are they different

jimthompson5910 (jim_thompson5910):

have you seen the identity \[\Large \cos^2(u) + \sin^2(u) = 1\] ??

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so because sin(u) = 4/5, we can square both sides to get sin^2(u) = 16/25 which means \[\Large \cos^2(u) + \sin^2(u) = 1\] \[\Large \cos^2(u) + \frac{16}{25} = 1\] now isolate cos(u)

OpenStudy (anonymous):

cosu= 3/5 ??

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

now if cos(v)=7/25, then what is sin(v) ?

OpenStudy (anonymous):

24/25 ???

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so we have these 4 facts sin(u) = 4/5 cos(u) = 3/5 cos(v) = 7/25 sin(v) = 24/25

jimthompson5910 (jim_thompson5910):

you plug them into the right side of cos(u+v) = cos(u)cos(v) - sin(u)sin(v) to compute cos(u+v)

OpenStudy (anonymous):

Is the answer A) -3/5 ????

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

oh wait, one sec

jimthompson5910 (jim_thompson5910):

yes sorry, I had the wrong values in my calculator

jimthompson5910 (jim_thompson5910):

cos(u+v) = -3/5

OpenStudy (anonymous):

great thank you

jimthompson5910 (jim_thompson5910):

you're welcome